OpenStudy (anonymous):

Let f be the function given by f(x)=3x^4 +x^3 -21x^2

OpenStudy (anonymous):

a.) Write an equation of the line tangent to the graph of f at the point (2,-28) b.) Find the absolute minimum value of f. Show the analysis that leads to your conclusion c.) Find the x-coordinate of each point of inflection on the graph of f. Show the analysis that leads to your conclusions

OpenStudy (anonymous):

@phi

OpenStudy (phi):

this looks like calculus for a) find the derivative of your curve, and sub in x=2 to find the slope at that point. once you have the slope, use the point slope formula for the equation of the tangent line.

OpenStudy (anonymous):

kay i got y=24x-76

OpenStudy (anonymous):

How about for part b.) and c.)

OpenStudy (anonymous):

@phi

OpenStudy (phi):

ok on a) for the min, set the derivative = 0 and solve for x see http://www.khanacademy.org/math/calculus/derivative_applications/critical_points_graphing/v/minima--maxima-and-critical-points it is short and helpful

OpenStudy (anonymous):

@phi so my answer for a.) y=24x-76 is wrong?

OpenStudy (anonymous):

or did u mean b.) not a.)

OpenStudy (phi):

I said you were ok on a). That means your answer for a is correct. the video is to help you finish the problem.

OpenStudy (anonymous):

would it be easier if i divided the derivative by 3: 12x^3 +3x^2 - 42x=0?

OpenStudy (phi):

of course, also factor out an x (x=0 is one solution)

OpenStudy (anonymous):

so ishould factor out an x and 3 o: 3x(4x^2 +x - 14)?

OpenStudy (phi):

3x(4x^2 +x - 14)=0

OpenStudy (anonymous):

ok so i got 3x=0 (4x-7)(1x+2)=0 x=0, 7/4, 2

OpenStudy (anonymous):

@phi

OpenStudy (phi):

you mean x=-2

OpenStudy (phi):

time to watch the video, and maybe the one after it where he does this same type of problem

OpenStudy (anonymous):

oh yeah sorry i mean -2 :)

OpenStudy (anonymous):

sorry i can't really watch the vid cause my cp is lagging x-x

OpenStudy (anonymous):

so would x=-2 be the absolute minimum value of f?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

ok how about for part c.)

OpenStudy (anonymous):

@agent0smith can you help me with part c.) please? :)

OpenStudy (agent0smith):

Part c just wants you to find every point where the gradient is zero. So, wherever f'(x) is equal to zero.

OpenStudy (agent0smith):

Which...it looks like you already did. x=0, 7/4, -2

OpenStudy (anonymous):

ok thank you can u help mewith another please ^^